# Anti-rotors and algebra invariants

### v35

@@ Abstract.
We derive previously unrecognized isomorphism invariants for real finite-dimensional unital associative algebras, and explore various implications regarding analysis of algebra structure.
@@ 1. Introduction
As the means of understanding algebra structure, identification of the invariants of an algebra is a task of fundamental importance as well as an open-ended endeavor given that the algebra classification problem is apparently intractable [1, 2]. While classical algebra isomorphism invariants include an algebra’s dimension, types of ideals, and invariants stemming from composition series (the Jordan-Hölder Theorem for modules), more recent examples are derived from (co)homology formalisms [3, 4, 5]. The invariants we present here belong to a different class.
@@ 2. The Skewer and the anti-rotor
Our program is motivated by examination of functions implied by the bilinear product operation between elements of an algebra $A$, as a means of studying the algebra itself. Each such unary operation, $f:D_{A,f}\rightarrow A$, is evidently a vector field (e.g., $f(s)=s^{-1}$, $f(s)=s^{2}$, etc.), where $D_{A,f}$ is an appropriate domain in the algebra’s vector space of elements. In examination of a vector field it is of interest to consider the associated infinitesimal rotation tensor (the exterior derivative of the differential 1-form dual to the vector field; its Hodge dual is the curl of the vector field). Its computation requires choice of a metric (i.e., deriving from a symmetric bilinear form, and also referred to as a scalar product or metric tensor). Specification of the standard topology entails imposition of the Euclidean inner product as the metric, allowing the computation to proceed. However, this choice of metric is arbitrary (i.e., not inherent in the axioms of an algebra), and thus can only be considered a default against which the results of all other choices should be compared. Hence, with respect to a specific choice of unary operation $f$ to study, we are led to consideration of the mapping,
(2.1) $\not{\text{S}}_{f}:Sym_{n}\rightarrow Skew_{n},$
@@
In (2.1), $Sym_{n}$ is the linear space of all possible metric tensors relevant to $\mathbb{R}^{n}$ (inclusive of metrics that are not necessarily positive-definite and not necessarily nondegenerate), $Skew_{n}$ is the linear space of all skew-symmetric tensors relevant to $\mathbb{R}^{n}$ (which is also the space of possible infinitesimal rotation tensors). Equation (2.2) is interpreted as follows. An algebra element is understood to be a column vector $s=(x_{1},\dots,x_{n})^{T}$, we define $\mathbf{d}s\equiv(dx_{1},\dots,dx_{n})^{T}$, and $\mu$ is a real symmetric matrix. On the right-hand-side of (2.2), $d$ represents the exterior derivative operation, thus acting on $(f(s))^{T}\mu\,\mathbf{d}s=\big{(}\mu f(s)\big{)}\cdot\mathbf{d}s$. Consequently, the right-hand-side of (2.2) yields the infinitesimal rotation tensor associated with the vector field $f(s)$ with respect to the metric $\mu$ - or, equivalently, the infinitesimal rotation tensor of vector field $\mu f(s)$ with respect to the Euclidean metric.
@@
$\not{\text{S}}_{f}$ is a linear transformation mapping symmetric tensors to skew-symmetric tensors, so we refer to $\not{\text{S}}_{f}$ as a “Skewer” (the slash through “symmetric” S represents a skewer, as with a shish-kabob). Since we assume the standard basis, $Sym_{n}$ can be represented by a parametrized symmetric $(n\times n)$-matrix with $\frac{n(n+1)}{2}$ independent parameters, i.e., it is given as a symmetric matrix with a different real parameter in each of its upper triangular entries. Every possible metric tensor is given by some realization of the set of parameters (a realization being the result of any particular choice of $\frac{n(n+1)}{2}$ real values for the members of the set of parameters). Consequently, this parametrized symmetric matrix is mapped by a Skewer to a parametrized skew-symmetric matrix representing the resulting infinitesimal rotation tensors. Each entry not on the main diagonal of the latter matrix is a rational function of the $n$ components of the elements of algebra $A$, and the above $\frac{n(n+1)}{2}$ parameters appear as coefficients of constituent rational function terms which sum to the entry.
While the image of a Skewer is thereby complicated, the kernel of a Skewer is much less so since components of the elements of $A$ play no role. Given $f$, an algebra is thus associated with the kernel of the Skewer. In other words, we have a mapping from algebras to subspaces of the space of metric tensors. We will show that analysis of this mapping has much to say about algebra structure.
Since “kernel of the Skewer” is a somewhat unwieldy expression, that entity will be referred to as the “anti-rotor”, because its members annihilate the curl of the vector field. For definiteness, in this paper we choose $f(s)=s^{-1}$. In fact, as indicated in Section 8, it might be sufficient to consider only this function.
@@ 3. The anti-rotor and resulting invariants
Using the left regular representation, it is to establish the existence of an open ball centered at the algebra’s multiplicative identity that is devoid of non-units.
@@
where we have used Definition 3.1, the commutativity of a differential with a linear transformation, and $(Ks_{1})^{-1}=K(s_{1}^{-1})$ (since $K$ is an isomorphism). It follows from (3.3) that every uncurling metric of algebra $A_{2}$ is associated with an uncurling metric of $A_{1}$, i.e., from the last equality in (3.3) we have $K^{T}L_{2}K\in\mathfrak{u}_{A_{1}}$. The isomorphism hypothesis allows us to apply an analogous argument employing $K^{-1}$, indicating that for $L_{1}\in\mathfrak{u}_{A_{1}}$ we have the associated $(K^{-1})^{T}L_{1}K^{-1}\in\mathfrak{u}_{A_{2}}$. Furthermore, $K^{T}L_{2}K\in\mathfrak{u}_{A_{1}}$ implies $(K^{-1})^{T}[K^{T}L_{2}K]K^{-1}=(K^{T})^{-1}[K^{T}L_{2}K]K^{-1}=L_{2}$. Similarly, $(K^{-1})^{T}L_{1}K^{-1}\in\mathfrak{u}_{A_{2}}$ is such that $K^{T}[(K^{-1})^{T}L_{1}K^{-1}]K=L_{1}$. The theorem follows from the implied bijection. ∎
@@
Consequently, $\mathfrak{u}$ can be represented as a symmetric $(n\times n)$-matrix $M_{\mathfrak{u}}$ whose entries are particular linear combinations of $m$ independent parameters (where $m$ is the dimension of $\mathfrak{u}$), because each uncurling metric results from a single set of selected real values for those parameters. The determinant of $M_{\mathfrak{u}}$, written as ${\rm det}(M_{\mathfrak{u}})$, is a polynomial in some or all of the components of an ordered $m$-tuple of parameters, $(\alpha_{1},\dots,\alpha_{m})$. If $m$ is small we will denote the parameters using different Greek letters like $\alpha$, $\beta$, $\gamma$, etc. In the sequel, $\mathfrak{u}$ and $M_{\mathfrak{u}}$ will be used interchangeably depending on the context.
@@
    $V_{\mathbb{R}}({\rm det}(M_{\mathfrak{u}}))=\{(\alpha_{1},\dots,\alpha_{m})\in\mathbb{R}^{m}:{\rm det}(M_{\mathfrak{u}})=0\}.$
@@ 4.2. Special unital norms
Let $\eta^{-1}$ be the set consisting of the multiplicative inverses of all members of $\eta$. It is easily shown that $\eta^{-1}$ is an open neighborhood of $\mathbf{1}$, and $\eta$, $\eta^{-1}$ are diffeomorphic via the multiplicative inversion operation. We now consider a smooth path $\mathcal{P}$ from $\mathbf{1}$ to $s$ which is inside $\eta$. Let $\mathcal{P}^{-1}$ be the point set consisting of the multiplicative inverses of the members of $\mathcal{P}$. Clearly, $\mathcal{P}^{-1}$ is a continuous path contained in $\eta^{-1}$. We then have,
@@ 6. Basic functorial aspects and a couple of useful results
We define category P as follows. An object is a symmetric $(n\times n)$-matrix where $n$ is any chosen positive integer, and where each matrix entry is a linear combination of the components of an $m$-tuple of distinct real parameters, $m\geq 1$. Furthermore, this parametrized matrix has the feature that if all of the $m$ parameters are replaced by 0 except for one parameter being replaced by 1, and one forms the collection of all such matrices except that a different one of the $m$ parameters is replaced by $1$ in each case, then the resulting $m$ real matrices are linearly independent. Thus, for any object of P, the implied set of real matrices resulting from all realizations of the associated $m$-tuple of parameters is an $m$-dimensional vector space (a realization of the $m$-tuple of parameters is the $m$-tuple resulting from a choice of real values for the $m$ parameters). An object of P is considered to be the same as the latter vector space, each of whose members is a real matrix, and such an object can be represented by a matrix whose entries are linear combinations of a set of real parameters. There is a morphism pointing from object $P$ to object $P^{\prime}$ if there exists a real matrix $w$ such that $ww^{T}$ is nonsingular and $w^{T}Pw$ is a subspace of $P^{\prime}$. We could similarly notate the morphism as $w$, but it turns out that would cause confusion once we introduce the functor. Therefore, we will instead suggestively notate the morphism by the symbol “ $w^{T}[\cdot]w$ ”. The composition of morphism $\begin{CD}P@>{w^{T}[\cdot]w}>{}>P^{\prime}\end{CD}$ with morphism $\begin{CD}P^{\prime}@>{w^{\prime T}[\cdot]w^{\prime}}>{}>P^{\prime\prime}\end{CD}$ is the morphism with source $P$ and target $P^{\prime\prime}$ resulting from $w^{\prime}w$ (i.e., analogous to how a morphism from $P$ to $P^{\prime}$ results from real matrix $w$). This morphism exists since it is easy to show that $(w^{\prime}w)^{T}P(w^{\prime}w)$ is a subspace of $P^{\prime\prime}$ and $(w^{\prime}w)(w^{\prime}w)^{T}$ is nonsingular. In other words, given the way we are notating the morphisms,
(6.1) $(w^{\prime T}[\cdot]w^{\prime})\circ(w^{T}[\cdot]w)=(w^{\prime}w)^{T}[\cdot](w^{\prime}w).$
@@
Based on the prior paragraph, morphism $F(k_{1})=k_{1}^{T}[\cdot]k_{1}$ exists and morphism $F(k_{2})=k_{2}^{T}[\cdot]k_{2}$ exists. But from (6.1), we have $F(k_{2}\circ k_{1})=F(k_{1})\circ F(k_{2})$. The theorem then follows. ∎
@@ 7. Illustrations
Taking $\alpha_{1},\dots,\alpha_{m}$ to be the parameters appearing in $M_{\mathfrak{u}}$, the integral on the right-hand-side of (4.3) can be written as
(7.1) $\log\mathfrak{u}^{d}=\int_{\mathbf{1}}^{s}[M_{\mathfrak{u}}(t^{-1})]\cdot\mathbf{d}t=\sum_{q=1}^{m}\left(\alpha_{q}\int_{\mathbf{1}}^{s}[\mathfrak{u}_{q}t^{-1}]\cdot\mathbf{d}t\right).$
On the right-hand-side, $\mathfrak{u}_{q}$ is defined from $M_{\mathfrak{u}}$ by setting $\alpha_{q}=1$ with all other parameters set to zero. All integrals are path-independent, and thus can be evaluated on a linear path of integration from $\mathbf{1}$ to $s$ - on which the integrands are rational functions of one variable. Thereby, the integration in principle yields a sum of rational function terms, logarithm terms, and arctangent terms. Nevertheless, in dimensions greater than four a closed form expression for the indefinite integration (where $s$ is a variable) is not possible because it would depend on expressions for the roots of the polynomial in the denominator of the rational function integrand, which will depend on the indefinite components of $s$.
However, in dimensions less than five the function obtained through evaluation of the integrals on the right-hand-side of (4.16) is computable in closed form, and given by respective sums of rational function terms, logarithm terms, and arctangent terms each multiplied by one of the parameters $\alpha_{q}$ arising in $M_{\mathfrak{u}}$. But the rational function terms can be collected into a single rational function term involving some number $\tau_{A,{\rm rat}}$ of distinct parameters. Furthermore, the function’s logarithm terms can be collected into a single (real) logarithm involving some number $\tau_{A,{\rm log}}$ of distinct parameters. Similarly, its arctangent terms, which can be alternatively expressed in the format $\text{arctan}(w)=\frac{i}{2}\log\frac{1-iw}{1+iw}$, can be collected into a single such logarithm term utilizing imaginary coefficients, involving some number $\tau_{A,{\rm arc}}$ of distinct parameters. Given two isomorphic algebras $A_{1}$, $A_{2}$, Theorem 4.4 implies that based on dimensional considerations relating to the respective numbers of distinct parameters associated with the single rational function term, the single real logarithm term, and the single logarithm term with imaginary coefficients, we must have

### v36

@@ Abstract.
We derive previously unrecognized geometrically-based isomorphism invariants for real finite-dimensional unital associative algebras, resulting from examination of transformations from an algebra to itself that are dependent on the algebra’s bilinear product operation.
@@ 1. Introduction
As the means of understanding algebra structure, identification of the invariants of an algebra is a task of fundamental importance as well as an open-ended endeavor given that the algebra classification problem is apparently intractable [1, 2]. While classical algebra isomorphism invariants include things like an algebra’s dimension, types of ideals, and invariants stemming from composition series (the Jordan-Hölder Theorem), more recently they have derived from module categories, e.g., involving homological formalisms such as found in [3, 4, 5]. The invariants we present here belong to a different class.
@@ 2. The Skewer and the anti-rotor
Our program begins with consideration of the transformations of an algebra $A$ to itself that are dependent on its bilinear product operation, as a means of studying the algebra. Each such unary operation, $f:D_{A,f}\rightarrow A$, is a function as well as vector field (e.g., $f(s)=s^{-1}$, $f(s)=s^{2}$, etc.), where $D_{A,f}$ is an appropriate domain in the algebra’s vector space of elements. In examination of a vector field it is of interest to consider the associated infinitesimal rotation tensor (the exterior derivative of the differential 1-form dual to the vector field; its Hodge dual is the curl of the vector field). Its computation requires choice of a metric (i.e., deriving from a symmetric bilinear form, and also referred to as a scalar product or metric tensor). Specification of the standard topology entails imposition of the Euclidean inner product as the metric, allowing the computation to proceed. However, this choice of metric is arbitrary (i.e., not inherent in the axioms of an algebra), and thus can only be considered a default against which the results of all other choices should be compared. Hence, with respect to a specific choice of function $f$ to study, we are motivated to examine the mapping,
(2.1) $\not{\text{S}}_{f}:\text{Sym}_{n}\rightarrow\delta\text{Rot}_{n},$
@@
In (2.1), $\text{Sym}_{n}$ is the linear space of all possible metric tensors relevant to $\mathbb{R}^{n}$ (inclusive of metrics that are not necessarily positive-definite and not necessarily nondegenerate), and $\delta\text{Rot}_{n}$ is the linear space of all possible infinitesimal rotation tensors relevant to $\mathbb{R}^{n}$. Equation (2.2) is interpreted as follows. An algebra element is understood to be a column vector $s=(x_{1},\dots,x_{n})^{T}$, we define $\mathbf{d}s\equiv(dx_{1},\dots,dx_{n})^{T}$, and $\mu$ is a real symmetric $(n\times n)$-matrix. On the right-hand-side of (2.2), $d$ represents the exterior derivative operation, thus acting on the 1-form $(f(s))^{T}\mu\,\mathbf{d}s=\big(\mu f(s)\big)\cdot\mathbf{d}s$. Consequently, the right-hand-side of (2.2) yields the infinitesimal rotation tensor associated with the vector field $f(s)$ with respect to the metric $\mu$ - or, equivalently, the infinitesimal rotation tensor of vector field $\mu f(s)$ with respect to the Euclidean metric.
@@
$\not{\text{S}}_{f}$ is a linear transformation mapping symmetric tensors to skew-symmetric tensors, so we refer to $\not{\text{S}}_{f}$ as a “Skewer” (the slash through “symmetric” S represents a skewer, as with a shish-kabob). Since we assume the standard basis, $\text{Sym}_{n}$ can be represented by a parametrized symmetric $(n\times n)$-matrix with $\frac{n(n+1)}{2}$ independent parameters, i.e., it is given as a symmetric matrix with a different real parameter in each of its upper triangular entries. Every possible metric tensor is given by some realization of the set of parameters (a realization being the result of any particular choice of $\frac{n(n+1)}{2}$ real values for the members of the set of parameters). Consequently, this parametrized symmetric matrix is mapped by a Skewer to a parametrized skew-symmetric matrix representing the resulting infinitesimal rotation tensors. Each entry not on the main diagonal of the latter matrix is a rational function of the $n$ components of the elements of algebra $A$, and the above $\frac{n(n+1)}{2}$ parameters appear as coefficients of constituent rational function terms which sum to the entry.
While the image of a Skewer is thereby complicated, the kernel of a Skewer is much less so because the components of the elements of $A$ do not appear. Furthermore, as indicated in Section 8, there is suggestive computational evidence that the kernel of the Skewer is independent of the choice of $f$ over the class of functions that are dependent on the algebra’s product operation (note that for $f$ not dependent on the algebra product, the kernel of the Skewer is simply $\text{Sym}_{n}$ itself). Notwithstanding its pleasing simplicity, if the resulting independence conjecture is false it will only mean that the collection of invariants proceeding from this program is larger than we anticipate, thereby increasing its power. Either way, it will be seen that analysis of the mapping of algebras to Skewer kernels reveals previously hidden aspects of algebra structure.
We will henceforth adopt “anti-rotor” as a more descriptive term for the kernel of the Skewer, because each of the latter’s members annihilates the curl of vector field $f$. For definiteness, in this paper we will work with $f(s)=s^{-1}$, since this choice is associated with a productive “dual” of the anti-rotor.
@@ 3. The anti-rotor and resulting invariants
Using the left regular representation, the existence of an open ball centered at the algebra’s multiplicative identity that is devoid of non-units is easily established.
@@
where we have used Definition 3.1, the commutativity of a differential with a linear transformation, and $(Ks_{1})^{-1}=K(s_{1}^{-1})$ (since $K$ is an isomorphism). It follows from (3.3) that every uncurling metric of algebra $A_{2}$ is associated with an uncurling metric of $A_{1}$, i.e., from the last equality in (3.3) we have $K^{T}L_{2}K\in\mathfrak{u}_{A_{1}}$. Injectivity of the implied map from $\mathfrak{u}_{A_{2}}$ to $\mathfrak{u}_{A_{1}}$ is evident due to the invertibility of $K$. The isomorphism hypothesis allows us to apply an analogous argument employing $K^{-1}$, indicating that for $L_{1}\in\mathfrak{u}_{A_{1}}$ we have the associated $(K^{-1})^{T}L_{1}K^{-1}\in\mathfrak{u}_{A_{2}}$, implying an injective map from $\mathfrak{u}_{A_{1}}$ to $\mathfrak{u}_{A_{2}}$. The theorem then follows from the bijection implied by the Cantor-Schoeder-Bernstein Theorem. ∎
@@
Consequently, $\mathfrak{u}$ can be represented as a symmetric $(n\times n)$-matrix $M_{\mathfrak{u}}$ whose entries are particular linear combinations of $m$ independent parameters (where $m$ is the dimension of $\mathfrak{u}$), because each uncurling metric results from a single set of selected real values for those parameters. The determinant of $M_{\mathfrak{u}}$, written as ${\rm det}(M_{\mathfrak{u}})$, is a polynomial in some or all of the components of an ordered $m$-tuple of parameters, $(\sigma_{1},\dots,\sigma_{m})$. If $m$ is small we will often denote the parameters using different Greek letters like $\alpha$, $\beta$, $\gamma$, etc. In the sequel, $\mathfrak{u}$ and $M_{\mathfrak{u}}$ will be used interchangeably depending on the context.
@@
    $V_{\mathbb{R}}({\rm det}(M_{\mathfrak{u}}))=\{(\sigma_{1},\dots,\sigma_{m})\in\mathbb{R}^{m}:{\rm det}(M_{\mathfrak{u}})=0\}.$
@@ 4.2. Special unital norms
Let $\eta^{-1}$ be the set consisting of the multiplicative inverses of all members of $\eta$. It is easily shown that $\eta^{-1}$ is an open neighborhood of $\mathbf{1}$, and $\eta$, $\eta^{-1}$ are diffeomorphic via the multiplicative inversion operation. We now consider a smooth path $\mathcal{P}$ from $\mathbf{1}$ to $s$ which is inside $\eta$. Let $\mathcal{P}^{-1}$ be the point set consisting of the multiplicative inverses of the members of $\mathcal{P}$. Clearly, $\mathcal{P}^{-1}$ is a smooth path contained in $\eta^{-1}$. We then have,
@@ 6. Basic functorial aspects and a couple of useful results
We define category P as follows. An object is a symmetric $(n\times n)$-matrix where $n$ is any chosen positive integer, and where each matrix entry is a linear combination of the components of an $m$-tuple of distinct real parameters, $m\geq 1$. Furthermore, this parametrized matrix has the feature that if all of the $m$ parameters are replaced by 0 except for one parameter being replaced by 1, and one forms the collection of all such matrices except that a different one of the $m$ parameters is replaced by $1$ in each case, then the resulting $m$ real matrices are linearly independent. Thus, for any object of P, the implied set of real matrices resulting from all realizations of the associated $m$-tuple of parameters is an $m$-dimensional vector space (a realization of the $m$-tuple of parameters is the $m$-tuple resulting from a choice of real values for the $m$ parameters). An object of P is considered to be the same as the latter vector space, each of whose members is a real matrix, and such an object can be represented by a matrix whose entries are linear combinations of a set of real parameters. There is a morphism pointing from object $P$ to object $P^{\prime}$ if there exists a real matrix $w$ such that $ww^{T}$ is nonsingular and $w^{T}Pw$ is a subspace of $P^{\prime}$. So we can say that the latter morphism results from $w$. We could similarly notate the morphism as $w$, but it turns out that would cause confusion once we introduce the functor. Therefore, we will instead suggestively notate the morphism implied by $w$ as the symbol “ $w^{T}[\cdot]w$ ”. The composition of morphism $\begin{CD}P@>{w^{T}[\cdot]w}>{}>P^{\prime}\end{CD}$ with morphism $\begin{CD}P^{\prime}@>{w^{\prime T}[\cdot]w^{\prime}}>{}>P^{\prime\prime}\end{CD}$ is the morphism with source $P$ and target $P^{\prime\prime}$ resulting from $ww^{\prime}$. This is truly a morphism since it is easy to verify $(ww^{\prime})^{T}P(ww^{\prime})\subset P^{\prime\prime}$ and $(ww^{\prime})(ww^{\prime})^{T}$ is nonsingular. Given the way we are notating the morphisms, we have
(6.1) $(w^{\prime T}[\cdot]w^{\prime})\circ(w^{T}[\cdot]w)=(ww^{\prime})^{T}[\cdot](ww^{\prime}).$
@@
Based on the prior paragraph, morphism $F(k_{1})=k_{1}^{T}[\cdot]k_{1}$ exists and morphism $F(k_{2})=k_{2}^{T}[\cdot]k_{2}$ exists. From observations made in the penultimate paragraph preceding this theorem’s statement, the composition of these morphisms then exists. According to the prior paragraph, $F(k_{2}\circ k_{1})$ also exists. Thus, from (6.1) and the definition of $F$, we have
$F(k_{2}\circ k_{1})=(k_{2}k_{1})^{T}[\cdot](k_{2}k_{1})=(k_{1}^{T}[\cdot]k_{1})\circ(k_{2}^{T}[\cdot]k_{2})=F(k_{1})\circ F(k_{2}).$
The theorem then follows. ∎
@@ 7. Illustrations
Taking $\sigma_{1},\dots,\sigma_{m}$ to be the parameters appearing in $M_{\mathfrak{u}}$, the integral on the right-hand-side of (4.3) can be written as
(7.1) $\log\mathfrak{u}^{d}=\int_{\mathbf{1}}^{s}[M_{\mathfrak{u}}(t^{-1})]\cdot\mathbf{d}t=\sum_{q=1}^{m}\left(\sigma_{q}\int_{\mathbf{1}}^{s}[\mathfrak{u}_{q}t^{-1}]\cdot\mathbf{d}t\right).$
On the right-hand-side, $\mathfrak{u}_{q}$ is defined from $M_{\mathfrak{u}}$ by setting $\sigma_{q}=1$ with all other parameters set to zero. All integrals are path-independent, and thus can be evaluated on a linear path of integration from $\mathbf{1}$ to $s$ - on which the integrands are rational functions of one variable. Thereby, the integration in principle yields a sum of rational function terms, logarithm terms, and arctangent terms. Nevertheless, in dimensions greater than four a closed form expression for the indefinite integration (where $s$ is a variable) is not possible because it would depend on expressions for the roots of the polynomial in the denominator of the rational function integrand, which will depend on the indefinite components of $s$.
However, in dimensions less than five the function obtained through evaluation of the integrals on the right-hand-side of (4.16) is computable in closed form, and given by respective sums of rational function terms, logarithm terms, and arctangent terms each multiplied by one of the parameters $\sigma_{q}$ arising in $M_{\mathfrak{u}}$. But the rational function terms can be collected into a single rational function term involving some number $\tau_{A,{\rm rat}}$ of distinct parameters. Furthermore, the function’s logarithm terms can be collected into a single (real) logarithm involving some number $\tau_{A,{\rm log}}$ of distinct parameters. Similarly, its arctangent terms, which can be alternatively expressed in the format $\text{arctan}(w)=\frac{i}{2}\log\frac{1-iw}{1+iw}$, can be collected into a single such logarithm term utilizing imaginary coefficients, involving some number $\tau_{A,{\rm arc}}$ of distinct parameters. Given two isomorphic algebras $A_{1}$, $A_{2}$, Theorem 4.4 implies that based on dimensional considerations relating to the respective numbers of distinct parameters associated with the single rational function term, the single real logarithm term, and the single logarithm term with imaginary coefficients, we must have